OpenStudy (anonymous):

f(x)=-2x-7 and g(x)=-4x+3 Find (f o g)(-5)

OpenStudy (anonymous):

f(g(x)) g(x) = -4x + 3 g(-5) = -4(-5) +3 g(-5) = -17 f(g(-5) f(-17) = -2x-7 f(-17) = -2(-17) - 7 f(-17) = 34 - 7 f(-17) = 27 f(g(-5) = 27

OpenStudy (anonymous):

nope im wrong, it should be 23 for g(-5) thus -56 -7 f(23) = -63 f(g(-5) = -63 always check ya work..

OpenStudy (anonymous):

thats wrong

OpenStudy (anonymous):

-53

OpenStudy (anonymous):

im just testing you

OpenStudy (anonymous):

yepp